OpenStudy (anonymous):

The numerator and the denominator of a fraction are thirteen and fifty, respectively. When a certain number is added to this numerator and subtracted from this denominator, the resulting fraction has a numerator to denominator ratio of two to one. What is the number?

OpenStudy (anonymous):

13 + x)/(50 - x) = 2 13 + x = 100 - 2x 3x = 87 x = 29